metallic rod 1 cm long with a square cross-section is
heated through 1°C. If Young's modulus of elasticity of
the metal is E and the mean coefficient of linear
expansion is alpha per degree Celsius, then the
compressional force required to prevent the rod from
expanding along its length is : (Neglect the change of cross
sectional area)

Grade:11

1 Answers

Aditya Gupta
3 years ago
 
Thermal expansion in form of change in length 
l2 = l1(1+alpha*t)
Young’s modulus = E 
E = (F/A)/(l/L)
E = FL/Al
L = l + del l 
E = F (l+del l )/A* del l 
Find the value of F 
F = EA(alpha *t)/(1+ alpha *t)
here, t= 1
so, F = EA(alpha)/(1+ alpha)
kindly approve :=)